# TOLEBI: Learning Fault-Tolerant Bipedal Locomotion via Online Status Estimation and Fallibility Rewards

### v1

@@ Abstract
With the growing employment of learning algorithms in robotic applications, research on reinforcement learning for bipedal locomotion has become a central topic for humanoid robotics. While recently published contributions achieve high success rates in locomotion tasks, scarce attention has been devoted to the development of methods that enable to handle hardware faults that may occur during the locomotion process. However, in real-world settings, environmental disturbances or sudden occurrences of hardware faults might yield severe consequences. To address these issues, this paper presents TOLEBI: A fault-tolerant learning framework for bipedal locomotion that handles faults on the robot during operation. Specifically, joint locking, power loss and external disturbances are injected in simulation to learn fault-tolerant locomotion strategies. In addition to transferring the learned policy to the real robot via sim-to-real transfer, an online joint status module incorporated. This module enables to classify joint conditions by referring to the actual observations at runtime under real-world conditions. The validation experiments conducted both in real-world and simulation with the humanoid robot TOCABI highlight the applicability of the proposed approach. To our knowledge, this work provides the first learning-based fault-tolerant framework for bipedal locomotion, thereby fostering the development of efficient learning methods in this field.
@@ I Introduction
Figure 1: TOLEBI, A framework for learning fault-tolerant bipedal locomotion. First, the motor failure events are injected during training in simulation to learn a fault-tolerant locomotion policy. Next, the policy is transferred to the real humanoid robot.
@@ II-B2 Learning-based Methods
To address these limitations, learning approaches have recently gained attention. As mentioned above, Reinforcement learning (RL) enables robots to learn robust control policies through interaction with diverse failure scenarios in simulation, eliminating the need for explicit fault modeling. In contrast, meta-learning techniques allow rapid adaptation to changing dynamics [1], while curriculum learning strategies gradually increase fault severity during training to improve robustness [17]. Recent works have proposed RL-based frameworks for quadrupedal robots that incorporate fault detection, recovery, and locomotion adaptation under degraded actuator conditions [12, 14, 28, 13]. Methods such as random joint masking [8] and multi-task learning [4] further enhance generalization across different failure modes.
@@
    First, a fault-tolerant learning framework for bipedal locomotion-TOLEBI is proposed, that incorporates a curriculum learning approach and motor failure simulations to facilitate sim-to-real transfer.
@@
    Second, we integrate the online joint status estimator, concurrently trained with the policy to infer joint status without additional training phases into TOLEBI.
@@
    Third, the fallibility reward is designed for TOLEBI under motor failure while preserving nominal locomotion style.
@@ III-A Reinforcement Learning
We formulate the bipedal locomotion control problem under motor failure as a Markov Decision Process (MDP). The MDP defined by a tuple ($\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma$), which is composed of the state space $\mathcal{S}$, the action space $\mathcal{A}$, a transition probability function $\mathcal{P}$, a reward function $\mathcal{R}$, and a discount factor $\gamma$. The agent aims to learn a policy $\pi_{\theta}(a|s)$ through interaction with the environment to maximize the expected discounted return $J(\pi_{\theta})$ defined as the cumulative discounted reward over a finite-horizon $T$:
@@ III-B Motor Failures
Motor failures in robots primarily arise from two sources. External disturbances, such as impacts, can mechanically constrain the actuators, while internal faults in electrical or control subsystems may disrupt power transmission or motor command signals. Both failure types can lead to significant performance degradation or even a complete loss of mobility during bipedal locomotion. This paper focuses on two failure scenarios: joint locking and power loss. The first, joint locking occurs when actuators become stuck, preventing any motion in the affected joints. The second, power loss, arises when joints remain physically free but cannot generate torque. Both failures severely compromise the robot’s balance, hinder the ability to track commanded velocities, and disrupt stable locomotion.
@@ IV TOLEBI - A Fault-Tolerant Learning Framework for Bipedal Locomotion
This section introduces TOLEBI, a reinforcement learning framework for fault-tolerant bipedal locomotion under unexpected motor failures. The proposed approach integrates motor failure simulation with joint masking, joint status estimation, curriculum learning and an action structure that incorporates phase modulation for adaptive gait timing. The entire control strategy is trained using Proximal Policy Optimization (PPO) [22] in Issac Gym and subsequently deployed on the real robot to ensure robustness operation.
@@ IV-A Overview
Our proposed framework, TOLEBI, illustrated in Fig. 2, integrates joint status estimation, observation history modeling, and curriculum learning to achieve fault-tolerant bipedal locomotion under motor failures. At each timestep, the observation includes proprioceptive observation, command velocities, and the estimated joint status, with a history buffer that stores the last $n_{history}=10$ entries with a stride of $n_{stride}=2$ between them. This approach allows the policy to leverage sequential information without a significant computational overhead. Motor failures are simulated by masking joint torque commands in Section IV-C, enabling the policy to learn control under partial actuation. The joint status estimator appends inferred joint status to the observation history, while curriculum learning in Section IV-D.
@@ IV-B2 Action Space
The action space $\mathcal{A}\in\mathbb{R}^{13}$ comprises 12 joint torque commands and an additional action for modulating the phase $\phi_{t}$. At each timestep, the policy outputs the mean $\mu_{\theta}$ of a Gaussian distribution $\mathcal{N}(\mu_{\theta},\sigma^{2})$, from which actions are sampled. The standard deviation $\sigma$ is predetermined according to the torque limits of individual joints for exploration.
The action space also includes a phase modulation action $a_{\delta\phi,t}$, which plays a critical role in adapting locomotion timing under motor failures. To enable rapid adaptation when actuators malfunction, the phase modulation action adjusts the motion period and timing as follows:
@@ IV-B3 Reward Function
Regulation Terms. Regulation terms are designed to improve motion smoothness, energy efficiency, and physical feasibility during locomotion. These terms penalize abrupt or excessive movements in body posture, joint dynamics, and interaction forces. The total regulation reward is formulated as follows:
@@
where each weight $w$ balances the contribution of the corresponding term.
@@ IV-C1 Failure Simulation
We simulate two types of motor failures—joint locking and power loss—by masking the action commands during training. For each training iteration, 90% of the environments are randomly assigned to an fault condition. Within this subset, the failure type is sampled with a 50% probability for each case. Then, a joint index $j\in\{0,1,\dots,11\}$ is uniformly selected, and its corresponding action is masked.
Joint Locking. The joint torque is computed using predefined proportional gains $K_{p}$ and derivative gains $K_{d}$ to hold the current joint position $q_{j}^{0}$ fixed at the moment of failure.
Power Loss. The commanded torque is set to zero, effectively disabling actuation.
Formally, the masked torque command $\tau_{j}$ for joint $j$ is given by
@@ IV-D Curriculum Learning
Curriculum learning plays an important role in stabilizing training for complex tasks like fault-tolerant humanoid locomotion. Rather than exposing the policy to all challenging conditions from the beginning, TOLEBI gradually increases task complexity based on the agent’s performance, allowing the policy to acquire fundamental locomotion skills before adapting to motor failures and external disturbances. As outlined in Algorithm 1, training begins with nominal locomotion under ideal conditions without any motor failures. This approach is taken because our preliminary experiments showed that premature exposure to faults led to unstable training and degraded performance. When the average episode length exceeds 20 seconds, motor failure simulation is introduced following the masking strategy described in Section IV-C, and the scale of the fallibility rewards is adjusted as detailed in Table I. where premature exposure to faults led to unstable training and degraded performance. Subsequently, when the average episode length under motor failures surpasses 24 seconds, push perturbations are applied to the robot’s base to improve robustness for sim-to-real transfer. This progressive curriculum provides a structured path from nominal locomotion to fault-tolerant and disturbance-resilient behaviors.
@@ V-A Experimental Setup
The proposed policy is trained using the PPO algorithm, with both the actor and critic networks modeled as two hidden layer MLPs consisting of 256 ReLU units per layer. Training is performed in the Isaac Gym simulator with 4096 parallel environments, running at a simulation frequency of 500 Hz and a control rate of 250 Hz. The maximum episode is set to 32 seconds. At each training iteration, a batch of 16,384 samples is collected, and the policy parameters are updated with a mini-batch size of 128 and a learning rate that linearly decays from $10^{-5}$ to $3\times 10^{-6}$.
@@ V-B Performance of Simulated Failure Scenarios
Table II reports the locomotion success rates across different motor failure scenarios in the Isaac Gym simulation. These rates are calculated by counting the number of successful episodes out of 4096 parallel environments, where a successful episode is defined as one that lasts for at least 20 seconds. We compare three methods: the baseline policy [6], the policy trained with joint masking and status estimation, and our full approach combining curriculum learning and the proposed fallibility reward. The results demonstrate the advantage of our method. Under joint locking conditions, the baseline policy fails to maintain stable locomotion in most cases, with multiple scenarios showing zero success rates. Incorporating joint masking and status estimation significantly improves performance, yet the success rate remains limited in challenging cases such as knee pitch and ankle pitch power loss failures. Our full approach achieves the highest average success rates under both joint locking (81.27%) and power loss (52.67%) conditions. The staged curriculum learning allows the policy to first acquire nominal locomotion skills before progressively handling motor failures and external perturbations, while the fallibility reward encourages robust recovery behaviors during fault-tolerance actuation. These results confirm that both components are essential for achieving fault-tolerant locomotion across diverse failure scenarios.
@@
Figure 3: Effect of the fallibility rewards. The reward mitigates early-contact impacts under motor failures and reduces impulsive forces that can reach up to 2000 N on the 100 kg humanoid robot TOCABI in real-world experiments.
@@ V-C Ablation Study
An ablation study was conducted to evaluate the contribution of the key components in TOLEBI: Joint status observation, fallibility rewards, phase modulation and curriculum learning. Table III presents the performance metrics—Root Mean Square Error (RMSE) and Mean Bias Error (MBE)—for both linear and angular velocity tracking averaged over motor failure scenarios. Removing the joint status observation led to a noticeable performance degradation, with particularly low performance under healthy conditions, indicating that online estimation of joint status is essential for accurate and fault-tolerant locomotion control. Excluding the fallibility rewards also reduced robustness under motor failures, as the policy failed to effectively promote safe behaviors, where the contact force reward specifically mitigates early-contact impacts that otherwise destabilize locomotion as also shown in Fig. 3. Similarly, removing the phase modulation resulted in the highest MBE among all ablation settings, as the absence of phase modulation failed to shorten the stance duration of the impaired leg, as this change eliminated the controller’s ability to adapt the gait timing. Furthermore, training without the curriculum learning prevented the policy from learning nominal locomotion style, resulting in gaits where the robot failed to properly lift its feet even under healthy conditions. The full model of TOLEBI, integrating all components, achieved the lowest tracking errors, highlighting their complementary benefits for robust fault-tolerant biped locomotion.

### v2

@@ Abstract
With the growing employment of learning algorithms in robotic applications, research on reinforcement learning for bipedal locomotion has become a central topic for humanoid robotics. While recently published contributions achieve high success rates in locomotion tasks, scarce attention has been devoted to the development of methods that enable to handle hardware faults that may occur during the locomotion process. However, in real-world settings, environmental disturbances or sudden occurrences of hardware faults might yield severe consequences. To address these issues, this paper presents TOLEBI: A fault-tolerant learning framework for bipedal locomotion that handles faults on the robot during operation. Specifically, joint locking, power loss and external disturbances are injected in simulation to learn fault-tolerant locomotion strategies. In addition to transferring the learned policy to the real robot via sim-to-real transfer, an online joint status estimator incorporated. This module enables to classify joint conditions by referring to the actual observations at runtime under real-world conditions. The validation experiments conducted both in real-world and simulation with the humanoid robot TOCABI highlight the applicability of the proposed approach. To our knowledge, this work provides the first learning-based fault-tolerant framework for bipedal locomotion, thereby fostering the development of efficient learning methods in this field.
@@ I Introduction
Figure 1: TOLEBI, A framework for learning fault-tolerant bipedal locomotion. Motor failures are injected during training in simulation to learn a fault-tolerant locomotion policy, and the policy is transferred to the real humanoid robot.
@@ II-B2 Learning-based Methods
To address these limitations, learning approaches have recently gained attention. As mentioned above, Reinforcement learning (RL) enables robots to learn robust control policies through interaction with diverse failure scenarios in simulation, eliminating the need for explicit fault modeling. In contrast, meta-learning techniques allow rapid adaptation to changing dynamics [1], while curriculum learning strategies gradually increase fault severity during training to improve robustness [16]. Recent works have proposed RL-based frameworks for quadrupedal robots that incorporate fault detection, recovery, and locomotion adaptation under degraded actuator conditions [13, 27, 12]. Methods such as random joint masking [8] and multi-task learning [4] further enhance generalization across different failure modes.
@@
    First, a fault-tolerant learning framework for bipedal locomotion TOLEBI is proposed that incorporates a curriculum learning approach and motor failure simulations to facilitate sim-to-real transfer.
@@
    Second, we integrate the online joint status estimator concurrently trained with the policy to infer joint status without additional training phases into TOLEBI.
@@
    Third, the fallibility reward is designed for TOLEBI under motor failure while preserving nominal locomotion.
@@ III-A Reinforcement Learning
We formulate the bipedal locomotion control problem under motor failure as a Markov Decision Process (MDP). The MDP defined by a tuple ($\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma$) that is composed of the state space $\mathcal{S}$, the action space $\mathcal{A}$, a transition probability function $\mathcal{P}$, a reward function $\mathcal{R}$, and a discount factor $\gamma$. The agent aims to learn a policy $\pi_{\theta}(a|s)$ by interacting with the environment to maximize the expected discounted return $J(\pi_{\theta})$ defined as the cumulative discounted reward over a finite-horizon $T$:
@@ III-B Motor Failures
Motor failures in robots primarily arise from two sources. External disturbances can mechanically constrain the actuators while internal faults in electrical or control subsystems might disrupt power transmission or motor command signals. Both failure types can lead to significant performance degradation or even a complete loss of mobility during bipedal locomotion. This paper focuses on two failure scenarios: Joint locking and power loss. Here, joint locking occurs when actuators become stuck and prevent any motion in the affected joints. In contrast, power loss arises when joints remain physically free but cannot generate torque. Both failures severely compromise the robot’s balance, hinder the ability to track commanded velocities, and disrupt stable locomotion.
@@ IV TOLEBI - A Fault-Tolerant Learning Framework for Bipedal Locomotion
This section introduces TOLEBI: A reinforcement learning framework for fault-tolerant bipedal locomotion under unexpected motor failures. The proposed approach integrates motor failure simulation with joint masking, joint status estimation, curriculum learning and an action structure that incorporates phase modulation for adaptive gait timing. The entire control strategy is trained using Proximal Policy Optimization (PPO) [21] in Issac Gym and deployed on the real robot to ensure robustness at operation.
@@ IV-A Overview
Our proposed framework TOLEBI shown in Fig. 2 integrates joint status estimation, observation history modeling, and curriculum learning for fault-tolerant bipedal locomotion under motor failures. At each time step, the observation includes proprioceptive observation, command velocities, and the estimated joint status with a history buffer that stores the latest $n_{history}=10$ entries with a stride of $n_{stride}=2$ between them. This approach allows the policy to leverage sequential information without significant computational overhead. Motor failures are simulated by masking joint torque commands in Section IV-C enabling the policy to learn control under partial actuation. The joint status estimator appends inferred joint status to the observation history during curriculum learning in Section IV-D.
@@ IV-B2 Action Space
The action space $\mathcal{A}\in\mathbb{R}^{13}$ comprises 12 joint torque commands and an additional action for modulating the phase $\phi_{t}$. At each timestep, the policy outputs the mean $\mu_{\theta}$ of a Gaussian distribution $\mathcal{N}(\mu_{\theta},\sigma^{2})$, from which actions are sampled. The standard deviation $\sigma$ is predetermined according to the torque limits of individual joints for exploration. The action space also includes a phase modulation action $a_{\delta\phi,t}$, which plays a critical role in adapting locomotion timing under motor failures. To enable rapid adaptation when actuators malfunction, the phase modulation action adjusts the motion period and timing as follows:
@@ IV-B3 Reward Function
Regulation Terms. We introduce the regulation terms to improve motion smoothness, energy efficiency, and physical feasibility during locomotion. These terms penalize sudden or excessive movements in body posture, joint dynamics, and interaction forces. The total regulation reward is given by
@@
where each weight $w$ balances the contribution of the respective term.
@@ IV-C1 Failure Simulation
We simulate two types of motor failures: Joint locking and power loss by masking the action commands during training. For each training iteration, 90% of the environments are randomly assigned to a fault condition. In this subset, the failure type is sampled with a 50% probability for each case. Next, a joint index $j\in\{0,1,\dots,11\}$ is uniformly selected to mask its corresponding action.
Joint Locking. The joint torque is computed via predefined proportional gains $K_{p}$ and derivative gains $K_{d}$ to hold the current joint position $q_{j}^{0}$ fixed at the moment of failure.
Power Loss. The commanded torque is set to zero to effectively disable actuation. Formally, the masked torque command $\tau_{j}$ for joint $j$ is given by
@@ IV-D Curriculum Learning
Curriculum learning plays an important role in stabilizing training for complex tasks like fault-tolerant humanoid locomotion. Rather than exposing the policy to all challenging conditions from the beginning, TOLEBI gradually increases task complexity based on the agent’s performance, allowing the policy to acquire fundamental locomotion skills before adapting to motor failures and external disturbances. As outlined in Algorithm 1 the training starts with nominal locomotion under ideal conditions without failures. This approach is taken because our preliminary experiments showed that premature exposure to faults led to unstable training and degraded performance. When the average episode length exceeds 20 seconds, motor failure simulation is introduced following the masking strategy in Section IV-C. The fallibility rewards are adjusted according to Table I. where premature exposure to faults led to unstable training and degraded performance. Subsequently, when the average episode length under motor failures surpasses 24 seconds, push perturbations are applied to the robot’s base to improve robustness for sim-to-real transfer. This progressive curriculum provides a structured path from nominal locomotion to fault-tolerant and disturbance-resilient behaviors.
@@ V-A Experimental Setup
The proposed policy is trained on the PPO algorithm with both the actor and critic networks modeled as two hidden layer MLPs consisting of 256 ReLU units per layer. The training is performed in the Isaac Gym simulator with 4096 parallel environments at a simulation frequency of 500 Hz and a control rate of 250 Hz. The maximum episode is set to 32 seconds. At each training iteration, a batch of 16,384 samples is collected, and the policy parameters are updated with a mini-batch size of 128 and a learning rate that linearly decays from $10^{-5}$ to $3\times 10^{-6}$.
@@ V-B Performance of Simulated Failure Scenarios
Table II reports the locomotion success rates across different motor failure scenarios in the Isaac Gym simulation. These rates are calculated by counting the number of successful episodes in 4096 parallel environments, where a successful episode is defined as one that lasts for at least 20 seconds. We compare three methods: The baseline policy [6], the policy trained with joint masking and status estimation, and our complete approach with the curriculum learning and the proposed fallibility reward. The results demonstrate the enhanced performance of our method. Under joint locking conditions, the baseline policy fails to maintain stable locomotion in most cases with multiple scenarios showing zero success rates. Incorporating joint masking and status estimation significantly improves the performance. The success rate remains limited in challenging cases like knee pitch and ankle pitch power loss failures. Our full approach achieves the highest average success rates under both joint locking (81.27%) and power loss (52.67%) conditions. The staged curriculum learning allows the policy to first acquire nominal locomotion skills prior to progressively handling motor failures and external perturbations while the fallibility reward encourages the robust recovery during fault-tolerance actuation. These results confirm that both components are essential for fault-tolerant locomotion across diverse failure scenarios.
@@
Figure 3: Effect of the fallibility rewards. The reward mitigates early-contact impacts under motor failures and reduces impulsive forces that can reach up to 2000 N on the 100 kg robot TOCABI in real-world experiments.
@@ V-C Ablation Study
This section evaluates TOLEBI’s key components: Joint status observation, fallibility rewards, phase modulation and curriculum learning. Table III shows the performance metrics Root Mean Square Error (RMSE) and Mean Bias Error (MBE) for both linear and angular velocity tracking averaged over motor failure scenarios. Removing the joint status observation led to a remarkable performance degradation with particularly low performance under healthy conditions. This indicates that estimating the joint status online is essential for accurate and fault-tolerant locomotion control. Excluding the fallibility rewards also reduced robustness under motor failures as the policy failed to effectively promote safe behaviors, where the contact force reward specifically mitigates early-contact impacts that otherwise destabilize locomotion as shown in Fig. 3. Similarly, removing the phase modulation resulted in the highest MBE among all ablation settings since the absence of phase modulation failed to shorten the stance duration of the impaired leg. This eliminated the controller’s ability to adapt the gait timing. Furthermore, training in the absence of curriculum learning prevented the policy from learning nominal locomotion that resulted in gaits where the robot failed to properly lift its feet even under healthy conditions. The complete model of TOLEBI with all components achieved the lowest tracking errors that highlights their benefits for robust fault-tolerant biped locomotion.